OpenStudy (anonymous):

One person can do a certain job in six days, a second person can do the same job in two days, and a third person can do this job in three days. How many days will they take to do the job together? 1 day 2 days 6 days

OpenStudy (anonymous):

\[\LARGE \frac{x}{6}+\frac x2+\frac x3=days \] \[\LARGE \frac{x}{6}+\frac {3x}{6}+\frac {2x}{6}=days \] \[\LARGE \frac{x+3x+2x}{6}=days \] \[\LARGE \frac{6x}{6}=days \] \[\LARGE x=days \] and this is weird !! O_O ... lol Although I think they all together can finish the job in 1 day...